Suppose the mean length of time between submission of a state tax return requesting a refund and the issuance of the refund is 47 days, with standard deviation 6 days. Find the probability that in a sample 50 of returns requesting a refund, the mean such time will be more than 50 days


1
Expert's answer
2021-11-26T13:24:57-0500

"\mu = 47 \\
\sigma = 6 \\
n=50 \\
P(\bar{X} > 50) = 1 -P(\bar{X} < 50) \\
= 1 -P(Z< \frac{50-47}{6 / \sqrt{50}}) \\
= 1 -P(Z< \frac{3}{0.848}) \\
= 1 -P(Z< 3.537) \\
= 1 -0.99979 \\
= 0.00021"